OpenStudy (anonymous):

(5-√11)(5-√11) 1) 36 2) 14 3) 25-10√11 4) 25-√11

OpenStudy (anonymous):

The right answer isn't among the options.

OpenStudy (anonymous):

25 + 11 - 2*5√11 36 - 10√11 the right ans is 36 - 10√11

OpenStudy (anonymous):

So would i be better off guessing 1 or 3?

OpenStudy (anonymous):

25-2(5*sqrt11)+11, 25-10*sqrt(11)+11, 36-10*sqrt(11) There's the answer.

OpenStudy (anonymous):

Are you sure you didn't mean: (5-√11)(5+√11)

OpenStudy (anonymous):

Wow they did

OpenStudy (anonymous):

Im sorry

OpenStudy (anonymous):

Use this: (a-b)(a+b)=a^2-b^2

OpenStudy (anonymous):

if it was (5-√11)(5+√11) = (5)² - (√11)² = 25 - 11 = 14 so 2) 14 is the right ans

OpenStudy (anonymous):

Thank you both